OpenStudy (anonymous):

Hello can someone help me solve this. x^2+y^2=74 y-x=2

OpenStudy (anonymous):

second equation tells you y = x+2 substitute in first equation to get the answer

OpenStudy (anonymous):

\[x^2+(x-2)^2=74\] \[x^2+x^2-4x+4=74\] \[2x^2-4x-70=0\]

OpenStudy (anonymous):

\[2(x-7)(x+5)=0\] \[x-7=0\] \[x=7\] or \[x+5=0\] \[x=5\]

OpenStudy (anonymous):

oops second should be x = -5

OpenStudy (anonymous):

oh damn. it was y - x = 2 so y = x+2

OpenStudy (anonymous):

so i was wrong. but doesn't change much except the -4x will be a +4x

OpenStudy (anonymous):

\[x^2+(x+2)^2=74\] \[x^2+x^2+4x+4=74\] \[2x^2+4x-70=0\] \[2(x-5)(x+7)=0\] \[x=5\] \[x=-7\]

OpenStudy (anonymous):

so y = 7 or y = -5 solutions are (5,7) and (-7,5)